# A Novel Benchmark for Few-Shot Semantic Segmentation in the Era of Foundation Models

### v1

@@ Abstract
In recent years, the rapid evolution of computer vision has seen the emergence of various vision foundation models, each tailored to specific data types and tasks. While large language models often share a common pretext task, the diversity in vision foundation models arises from their varying training objectives. In this study, we delve into the quest for identifying the most effective vision foundation models for few-shot semantic segmentation, a critical task in computer vision. Specifically, we conduct a comprehensive comparative analysis of four prominent foundation models: DINO V2, Segment Anything, CLIP, Masked AutoEncoders, and a straightforward ResNet50 pre-trained on the COCO dataset. Our investigation focuses on their adaptability to new semantic segmentation tasks, leveraging only a limited number of segmented images. Our experimental findings reveal that DINO V2 consistently outperforms the other considered foundation models across a diverse range of datasets and adaptation methods. This outcome underscores DINO V2’s superior capability to adapt to semantic segmentation tasks compared to its counterparts. Furthermore, our observations indicate that various adapter methods exhibit similar performance, emphasizing the paramount importance of selecting a robust feature extractor over the intricacies of the adaptation technique itself. This insight sheds light on the critical role of feature extraction in the context of few-shot semantic segmentation. This research not only contributes valuable insights into the comparative performance of vision foundation models in the realm of few-shot semantic segmentation but also highlights the significance of a robust feature extractor in this domain.
Index Terms—  Deep Learning, Foundation models, Semantic Segmentation, Few-Shot Learning
@@
Fig. 1: mIoU of various pretrained models when adapted to 3 different 1-shot segmentation tasks (only best performance achieved out of 5 tested methods is displayed).
1 Introduction
Semantic segmentation, the task of pixel-level object classification within images, i.e, assigning a class to every pixel in the image, plays a pivotal role in the landscape of computer vision applications [1]. In recent years, few-shot semantic segmentation has emerged as a challenging yet essential frontier [2]. This paradigm seeks to equip models with the ability to adapt to novel object classes with minimal labeled data. A standard pipeline for few-shot semantic segmentation typically comprises two key components: a pretrained feature extractor, often derived from a vision foundation model, and an adapter method. This setup leverages the knowledge embedded within the feature extractor and adapts it to the limited labeled data without the risk of overfitting. While extensive discussions regarding the selection of feature extractors for few-shot classification have been well-documented [3], the question of the optimal feature extractor for few-shot semantic segmentation remains largely unaddressed in the literature, to the best of our knowledge.
The primary objective of our research is to investigate which vision foundation model yields the most effective few-shot semantic segmentation pipeline. To this end, we introduce a novel benchmark, constructed using three widely-recognized semantic segmentation datasets, and systematically evaluate multiple adaptation methods. The four foundation models under consideration are DINO V2 [4], Segment Anything (SAM) [5], CLIP [6], and a Masked AutoEncoder [7] pretrained on ImageNet-1K. Our experimental results consistently demonstrate the superior performance of DINO V2 across various datasets.
In summary, our study makes the following key contributions:
@@ 1 Introduction
Our findings indicate that DINO V2 consistently outperforms all other models across multiple datasets and adaptation methods, with SAM yielding competitive results, particularly on specific datasets. Additionally, we find that simple and efficient adaptation methods produce results comparable to more complex counterparts. We anticipate that our research will assist newcomers in selecting appropriate solutions for their few-shot semantic segmentation tasks and contribute to the rigorous benchmarking of this evolving field.
@@ 2.1 Task Formulation
where $n$ represents the total number of classes or objects, $\mathrm{TP}_{i}$ is the true positive count for class $i$, $\mathrm{FP}_{i}$ is the false positive count for class $i$, and $\mathrm{FN}_{i}$ is the false negative count for class $i$. One noteworthy distinction from few-shot classification tasks is that individual images typically contain instances from various classes of interest. As such, it is challenging to define tasks in the traditional $k$-shot manner, where each class is associated with exactly $k$ elements in the support set.
In our proposed framework, we introduce a more practical approach. We define a $k$-shot segmentation task as one in which the support set $\mathcal{S}$ contains at least $k$ instances of each class. In a scenario involving a total of $n$ classes of interest, a $k$-shot task mandates the inclusion of precisely $nk$ unique images within the support set.
@@ 2.2 Benchmark Generation
    For each class $i$ (excluding the background class) within the chosen training dataset $\mathcal{D}$, we compile a list of all training images that contain at least one instance of that class. This results in $\mathcal{D}_{i}=\{\text{image}\in\mathcal{D}\mid\text{image contains instancesof class }i\}$.
@@
    Following an arbitrary order of classes, we randomly select $k$ images from $\mathcal{D}_{i}$ without replacement to be incorporated into the support set $\mathcal{S}$. During this process, we ensure that images previously selected are excluded from consideration.
@@
    In the rare event that there are insufficient remaining images for a particular class $i$, we reset the support set to empty $\mathcal{S}=\emptyset$ and recommence the procedure from Step 1. It is important to note that this reset condition did not occur in the course of our experiments.
Fig. 2: The overall pipeline of the Linear method: we freeze the backbone and train the shallow decoder on support set, and then we test the encoder+decoder on the query set
2.3 Datasets
In constructing our proposed benchmark, we have selected three prominent semantic segmentation datasets, each offering unique characteristics and challenges:
@@ 2.3 Datasets
PPD (PLANT PHENOTYPING DATASETS) [11] is a small dataset used for multiple tasks including plant segmentation (foreground to background segmentation) and leaf segmentation (multi-instance segmentation). We focus in our work on the plant segmentation task, where the dataset contains 783 images, 2 classes (foreground and background). Typical input resolutions for PPD images hover around 500 pixels for both width and height.
@@ 3 Backbones and Adaptation Methods
In the upcoming subsections, we provide comprehensive details on the specific models and adaptation methods that constitute the experimental foundation of our study.
@@ 3.1 Pretrained Models
In our study, we investigate five pretrained models, each offering unique characteristics and advantages:
@@
MAE (Masked AutoEncoder) [7] is pretrained on ImageNet-1K [13] using a self-supervised technique involving masking portions of input images and reconstructing them using partial tokens from the unmasked segments. It comprises a Vision Transformer (ViT) [14] component that maps observed image regions into latent representations and a decoder for image reconstruction. Previous research [7] has demonstrated MAE’s effectiveness in transfer learning for segmentation tasks, particularly in semantic segmentation.
@@
SAM (Segment Anything Model) [5] is a recent foundation model for segmentation, pretrained on the SA-1B dataset, which includes 11 million images and 1 billion segmentation masks. SAM exhibits robust generalization capabilities across a broad range of segmentation tasks. It utilizes a Vision Transformer (ViT) [14] initialized with MAE [7], a prompt encoder, and a mask decoder. SAM’s image encoder generates embeddings from input images, which are subsequently employed by the mask decoder in conjunction with the prompt encoder. While SAM’s mask decoder outputs class-agnostic masks intended for segmenting instances or objects at various granularities, our specific task focuses exclusively on the encoder’s role as a feature extractor and does not require prompts.
@@
DINO v2 [4] is a foundational vision model trained using the self-supervised DINO technique [15]. The model comprises a ViT trained on a diverse range of datasets, including notable semantic segmentation datasets like Cityscapes [9], ADE20k [16, 17], and Pascal VOC 2012 [18]. DINO v2 has previously been evaluated on semantic segmentation datasets in [4], utilizing the frozen backbone and a simple segmentation decoder.
@@
CLIP (Contrastive Language-Image Pre-training) [6] is a model that consists of two encoders, a vision encoder, and a text encoder. These encoders are jointly trained using a contrastive loss to align embeddings of images with their corresponding captions. The model is pretrained on a private dataset of 400 million (image, text) pairs and has gained significant attention for its utility in zero-shot classification on downstream tasks. In our study, we exclusively leverage ViT version of the visual encoder component of CLIP.
@@
FCN (Fully Convolutional Networks) [19] represents a conventional convolutional architecture extensively used for semantic segmentation tasks. It comprises an encoder and a decoder, with the former composed of convolutional layers and the latter incorporating deconvolutional layers. In our configuration, the encoder is based on ResNet50 [20] and is pretrained on a subset of COCO, specifically using the 20 classes shared with the Pascal VOC dataset. We incorporate FCN as a convolution-only benchmark for comparison.
@@ 3.2 Methods
We then consider five adaptation methods in our study:
Linear: A common approach [4] consists in freezing the pretrained backbone and adding a segmentation head, called decoder, consisting of a batch normalization layer followed by a convolutional 1x1 layer. Figure 2 illustrates the overall pipeline for this linear adaptation method.
Multilayer: It has been shown in [4] that using a concatenation of multiple layers from the pretrained model as output can boost performance in downstream tasks. We thus propose to use a linear method on the concatenation of the last 4 blocks of the pretrained model.
Singular Value Fine-tuning [21] (SVF): The SVF method decomposes parameters into three matrices using Singular Value Decomposition (SVD) and fine-tunes only the singular values while keeping the rest frozen. In [21], this approach demonstrated promising results in few-shot semantic segmentation tasks, showcasing its ability to specialize the pretrained model without overfitting. In our study, we initially apply the linear method and then proceed to fine-tune the singular values alongside the decoder parameters.
Low-Rank Adaptation [22] (LoRA): In this approach, we maintain the pretrained backbone’s frozen state and introduce low-rank trainable adapters, as described in [22]. Similar to the SVF method, LoRA enables tuning only a limited number of parameters. Our procedure begins with training the decoder using the linear method and then introduces LoRA adapters to the encoder. Subsequently, we fine-tune both the adapters and decoder parameters.
@@ 4.1 Training procedure
To ensure the reproducibility of our results, we conducted all experiments with three fixed random seeds and computed the mean Intersection-over-Union (mIoU) across these runs. For preprocessing, we followed the approach outlined in [23]. This involved random horizontal flipping, random scaling of the shorter side within the range of [400, 1600] while preserving the aspect ratio, and random cropping to a size of 1024x1024. We applied a subset of RandAug [24] operations, including auto contrast, equalize, rotate, color, contrast, brightness, and sharpness. Subsequently, all images were resized to 1024x1024 resolution, except for CLIP and MAE, which accepted input sizes of 224x224. We acknowledge that the resolution variation, depending on the considered model, may introduce some analysis challenges but was an unavoidable aspect of our setup. For the learning rate, we employed the poly learning rate scheduler with a power of 0.9 for the linear and multilayer methods. Specifically, we used a learning rate of 0.2 for Cityscapes, 0.05 for COCO, and 0.001 for PPD. For fine-tuning methods (SVF, LoRA, and fine-tune) we did a grid search across models and datasets with values including $10^{-2}$, $10^{-3}$, $10^{-4}$, $10^{-5}$, and $10^{-6}$ to determine the one yielding the best average performance. Linear and multilayer methods were trained with a batch size of 4. For the Fine-tuning, SVF, and LoRA methods, we initiated the training with the linear method and subsequently performed fine-tuning with the reduced batch size of 2, addressing memory limitations. Training duration extended to 200 epochs for Cityscapes and PPD, and 100 epochs for COCO. The query set consisted of the standard validation sets for Cityscapes and COCO. For PPD, given the absence of a conventional split, we performed a fixed 50/50 random split into training and validation sets. Our models were trained on a combination of NVIDIA A6000 and NVIDIA RTX3090 GPUs. Runtime varied across datasets and models, ranging from approximately 30 minutes per run for MAE, CLIP, and ResNet on Cityscapes to approximately 4 hours per run for DINO V2 on COCO. For the size of the models, we choose the base version of the vision transformers, leading to a 89M parameter model for SAM and 86M for DINO V2, CLIP and MAE. For the FCN we chose a ResNet50 as an encoder, leading to a 23M parameter model.
@@
Table 1: mIoU averaged over our three considered datasets in 1-shot few-shot semantic segmentation.

### v2

@@ Abstract
In recent years, the rapid evolution of computer vision has seen the emergence of various foundation models, each tailored to specific data types and tasks. In this study, we explore the adaptation of these models for few-shot semantic segmentation. Specifically, we conduct a comprehensive comparative analysis of four prominent foundation models: DINO V2, Segment Anything, CLIP, Masked AutoEncoders, and of a straightforward ResNet50 pre-trained on the COCO dataset. We also include 5 adaptation methods, ranging from linear probing to fine tuning. Our findings show that DINO V2 outperforms other models by a large margin, across various datasets and adaptation methods. On the other hand, adaptation methods provide little discrepancy in the obtained results, suggesting that a simple linear probing can compete with advanced, more computationally intensive, alternatives  [^1] Code to be released at: https://github.com/RedaBensaidDS/Foundation_FewShot
Keywords:
Deep Learning Foundation models Semantic Segmentation Few-Shot Learning
1 Introduction
Semantic segmentation, the task of pixel-level object classification within images, plays a pivotal role in computer vision applications [22]. In recent years, few-shot semantic segmentation has emerged as a challenging yet essential frontier [31]. This paradigm seeks to adapt models to novel object classes with minimal labeled data. A standard pipeline for few-shot semantic segmentation typically comprises two key components: a pretrained feature extractor, often derived from a vision foundation model, and an adapter method [2]. This setup leverages the knowledge embedded within the feature extractor and adapts it to the limited labeled data without the risk of overfitting. While extensive discussions regarding the selection of feature extractors for few-shot classification have been well-documented [21], the question of the optimal feature extractor for few-shot semantic segmentation remains largely unaddressed in the literature.
@@ 1 Introduction
Figure 1: mIoU of various pretrained models when adapted to 3 different 1-shot segmentation tasks (only best performance achieved out of 5 tested methods is displayed).
The primary objective of our research is to investigate which vision foundation model yields the most effective few-shot semantic segmentation pipeline. We are the first, to our knowledge, to systematically test the adaptation capabilities of renowned foundation models to this specific challenge. To this end, we introduce a novel benchmark, constructed using three widely-recognized semantic segmentation datasets, and systematically evaluate multiple adaptation methods. The four foundation models under consideration are DINO V2 [27], Segment Anything (SAM) [15], CLIP [28], and a Masked AutoEncoder [12] pretrained on ImageNet-1K. Our experimental results consistently demonstrate the superior performance of DINO V2 across various datasets.
In summary, from our study we claim the following key contributions:
@@
  • We conduct extensive experiments to better understand the elementary contribution of foundation models design (e.g. type of architecture, size, training dataset) onto the obtained performance.
Our findings indicate that DINO V2 consistently outperforms all other models across multiple datasets and adaptation methods, with SAM yielding competitive results, particularly on some datasets (Figure 1). Additionally, we find that simple and efficient adaptation methods produce results comparable to more complex counterparts, showing yet again the importance of feature extraction on top of that of downstream segmentation heads. We anticipate that our research will assist in selecting appropriate solutions for few-shot semantic segmentation tasks and contribute to the rigorous benchmarking of this evolving field.
@@ 2.1 Task Formulation
where $n$ represents the total number of classes or objects, $\mathrm{TP}_{i}$ is the true positive count for class $i$, $\mathrm{FP}_{i}$ is the false positive count for class $i$, and $\mathrm{FN}_{i}$ is the false negative count for class $i$. One noteworthy distinction from few-shot classification is that individual images typically contain instances from various classes of interest. As such, it is challenging to define tasks in the $k$-shot manner, where each class is associated with exactly $k$ elements in the support set.
In our proposed framework, we introduce a more practical approach. We define a $k$-shot segmentation task as one in which the support set $\mathcal{S}$ contains at least $k$ instances of each class. Using our proposed sampling protocol described in the next subsection, in a scenario involving a total of $n$ classes, a $k$-shot task mandates the inclusion of precisely $nk$ unique images within the support set.
@@ 2.2 Benchmark Generation
    For each class $i$ (excl. the background class) within the chosen dataset $\mathcal{D}$, we compile a list of all training images that contain at least one instance of that class, resulting in $\mathcal{D}_{i}=\{\text{image}\in\mathcal{D}\mid\text{image contains instancesof class }i\}$.
@@
    Following an arbitrary order of classes, we randomly select $k$ images from $\mathcal{D}_{i}$ without replacement to be incorporated into the support set $\mathcal{S}$. We ensure that images are only selected once.
@@
    In the event that there are insufficient remaining images for a particular class $i$, we restart from Step 1. It is important to note that this reset condition did not occur in the course of our experiments.
3 Datasets
In constructing our proposed benchmark, we have selected three prominent semantic segmentation datasets, each offering unique characteristics:
@@ 3 Datasets
PPD (PLANT PHENOTYPING DATASETS) [23] is a small dataset used for multiple tasks including plant segmentation (foreground to background) and leaf segmentation (multi-instance segmentation). We focus in our work on the plant segmentation task, where the dataset contains 783 images and 2 classes (foreground and background). Typical input resolutions for PPD images hover around 500 pixels for both width and height. Contrary to the two other datasets, we considered the background to be a class of its own for benchmark generation, meaning that a 1-shot problem would contain 2 support images.
@@ 4 Backbones and Adaptation Methods
Next, we provide comprehensive details on the specific models and adaptation methods that constitute the experimental foundation of our study.
@@ 4.1 Pretrained Models
In our study, we investigate five pretrained models, each offering unique characteristics and advantages (see also summary in Table 13):
@@
MAE (Masked AutoEncoder) [12] is a model pretrained on ImageNet-1K [6] using a self-supervised technique involving masking portions of input images and reconstructing them using partial tokens from the unmasked segments. It comprises a Vision Transformer (ViT) [7] component that maps observed image regions into latent representations and a decoder for image reconstruction. Previous research [12] has demonstrated MAE’s effectiveness in transfer learning for segmentation tasks, particularly in semantic segmentation.
@@
SAM (Segment Anything Model) [15] is a recent foundation model for segmentation, pretrained on the SA-1B dataset, which includes 11 million images and 1 billion segmentation masks. SAM exhibits robust generalization capabilities across a broad range of segmentation tasks. It utilizes a Vision Transformer (ViT) [7] initialized with a MAE [12], a prompt encoder, and a mask decoder. SAM’s image encoder generates embeddings from input images, which are subsequently employed by the mask decoder in conjunction with the prompt encoder. While SAM’s mask decoder outputs class-agnostic masks intended for segmenting instances or objects at various granularities, our specific task focuses exclusively on the encoder’s role as a feature extractor and does not require prompts.
@@
DINO v2 [27] is a foundational vision model trained using the self-supervised technique [1] called DINO. This methods consists in having two networks with the same architecture, namely the teacher and the student. The student parameters are learned by minimizing a loss measuring the discrepancy between teacher and student outputs learning a “local to global” correspondence where the student processes global and local views of the image while the teacher only processes the global views. The teacher is built with an exponential moving average of past iteration of the student. The model comprises a ViT trained on a diverse range of datasets, including notable semantic segmentation datasets like Cityscapes [3], ADE20k [42, 43], and Pascal VOC 2012 [8]. DINO v2 has previously been evaluated on (not few-shot) semantic segmentation datasets in [27], utilizing the frozen backbone and a simple segmentation decoder.
@@
CLIP (Contrastive Language-Image Pre-training) [28] is a model that consists of two encoders: a vision encoder and a text encoder. These encoders are jointly trained using a contrastive loss to align embeddings of images with their corresponding captions. The model is pretrained on a private dataset of 400 million (image, text) pairs and has gained significant attention for its utility in zero-shot classification on downstream tasks. In our study, we exclusively leverage the ViT version of the visual encoder component of CLIP.
@@
FCN (Fully Convolutional Network) [20] is a conventional convolutional architecture extensively used for semantic segmentation tasks. It comprises an encoder and a decoder, with the former composed of convolutional layers and the latter incorporating deconvolutional layers. In our configuration, the encoder is based on ResNet50 [13] and is pretrained on a subset of COCO, specifically using the 20 classes shared with the Pascal VOC dataset. We incorporate FCN as a convolution-only benchmark for comparison.
Figure 2: The pipeline of the Linear method: we freeze the backbone and train the shallow decoder on the support set, and then we test the encoder+decoder on the query set.
@@ 4.2 Methods
We then consider five adaptation methods in our study, ranging from linear probing methods to partial fine-tuning and full fine-tuning methods:
Linear: A common approach [27, 33] consists in freezing the pretrained backbone and adding a segmentation head, called decoder, consisting of a batch normalization layer followed by a convolutional 1x1 layer. Figure 2 illustrates the overall pipeline for this linear adaptation method.
Multilayer: It has been shown in [27] and [38] that using a concatenation of multiple layers from the pretrained model can boost performance in downstream tasks. We thus propose to use a linear method on the concatenation of the last 4 blocks of the pretrained model.
Singular Value Fine-tuning [34] (SVF): The SVF method decomposes parameters into three matrices using Singular Value Decomposition (SVD) and fine-tunes only the singular values diagonal matrix while keeping the rest frozen. In [34], this approach demonstrated promising results in few-shot semantic segmentation tasks, showcasing its ability to specialize the pretrained model without overfitting. In our study, we initially apply the linear method and then proceed to fine-tune the singular values of the encoder alongside the decoder parameters.
Low-Rank Adaptation [14] (LoRA): In this approach, we maintain the pretrained backbone’s frozen state and introduce low-rank trainable adapters, as described in [14]. Similarly to the SVF method, LoRA enables tuning only a limited number of parameters. Our procedure begins with training the decoder using the linear method and then introduces LoRA adapters to the encoder. Subsequently, we fine-tune both the adapters and decoder parameters.
@@ 5.1 Training Procedure
To ensure the reproducibility of our results, we conducted all experiments with three fixed random seeds and computed the mean Intersection-over-Union (mIoU) across these runs. For preprocessing, we followed the approach outlined in [10]. This involved random horizontal flipping, random scaling of the shorter side within the range of [400, 1600] while preserving the aspect ratio, and random cropping to a size of 1024x1024. We applied a subset of RandAug [4] operations, including auto contrast, equalize, rotate, color, contrast, brightness, and sharpness. Subsequently, all images were resized to 1024x1024 resolution, except for CLIP and MAE, which accept input sizes of 224x224. We acknowledge that the resolution variation, depending on the considered model, may introduce some analysis challenges but was an unavoidable aspect of our setup. For the learning rate, we employed the poly learning rate scheduler with a power of 0.9 for the linear and multilayer methods. Specifically, we used a learning rate of 0.2 for Cityscapes, 0.05 for COCO, and 0.001 for PPD. For fine-tuning methods (SVF, LoRA, and fine-tune), we did a grid search across models and datasets with values including $10^{-2}$, $10^{-3}$, $10^{-4}$, $10^{-5}$, and $10^{-6}$ to determine the one yielding the best average performance. Linear and multilayer methods were trained with a batch size of 4. For the Fine-tuning, SVF, and LoRA methods, we initiated the training with the linear method and subsequently performed fine-tuning with the reduced batch size of 2, addressing memory limitations. Training duration extended to 200 epochs for Cityscapes and PPD, and 100 epochs for COCO. The query set consisted of the standard validation sets for Cityscapes and COCO. For PPD, given the absence of a conventional split, we performed a fixed 50/50 random split into training and validation sets. Our models were trained on a combination of NVIDIA A6000 and NVIDIA RTX3090 GPUs. Runtime varied across datasets and models, ranging from approximately 30 minutes per run for MAE, CLIP, and ResNet on Cityscapes to approximately 4 hours per run for DINO V2 on COCO. For the size of the models, we chose the base version of the vision transformers, leading to a 89M parameters model for SAM and 86M for DINO V2, CLIP and MAE. For the FCN we chose a ResNet50 as an encoder, leading to a 23M parameters model.
5.2 What Works The Best?
First, we aim to identify the optimal combination of pretrained models and adaptation methods based on mIoU. Table 1 presents the average performance on the three considered datasets for all possible combinations of pretrained models and adaptation methods, along with their standard deviation, in the context of 1-shot problems. Several conclusions can be drawn:
  • SAM encoder exhibits strong performance on average. It outperforms CLIP, MAE, and ResNet on both Cityscapes and PPD. However, it yields comparatively poor results on COCO (refer to Appendix Table 9). This difference is likely attributed to our decision to retain the SAM image encoder while discarding the original mask decoder.
  • The multilayer method outperforms the linear approach for most models, indicating that including earlier feature maps, which are more locally focused [29, 7], provides the decoder with greater segmentation flexibility. Furthermore, it exhibits a lower standard deviation, indicating greater stability across the different runs of our proposed benchmark.
  • Despite variations in the number of trainable parameters (less than 1% for LoRA, around 2% for SVF, and 100% for finetune), the finetuning methods (SVF, LoRA, and Full Fine-tuning) yield comparable results between them.
@@ 5.2 What Works The Best?
Table 1: mIoU averaged over our three considered datasets in 1-shot few-shot semantic segmentation. Bold numbers correspond to the best performance across feature extractors and underline numbers to the best performance across adaptation methods.